Equal weights of methane and oxygen are mixed in empty container at 25 degree celcius.the fraction of total pressure exerted by oxygen is ..............

Grade:12

1 Answers

India VK
10 years ago
Ratio by weights = 1:1 Ratio by moles = 1/16 : 1/32 = 2:1 So fraction of pressure exerted by oxygen = mole fraction of oxygen = 1/3